# Towards Observation Lakehouses: Living, Interactive Archives of Software Behavior

### v1

@@ I Introduction
Most code-generating LLMs are trained on static artifacts – source code, comments, and specifications – rather than on materializations of run‑time behavior [NEURIPS2024_6efcc7fd]. Such models inevitably internalize patterns from buggy or mislabeled code as readily as from correct implementations, because they are rarely, if ever, exposed to trustworthy evidence of functional run-time behavior (i.e., functionality) [chen2021evaluatinglargelanguagemodels, 10174227]. Improving reliability, therefore, requires exposing models to ground‑truth behavior captured under execution.
@@
Rice’s theorem rules out deciding non‑trivial semantic properties of programs in general [riceTheorem], so for real systems the only practical way of doing this is through dynamic observation – that is, to execute software where it is expected to meet its advertised functionality and record what it actually does, either during controlled tests (i.e., software testing [ammann2017introduction]) or during real usage (i.e., monitoring and tracing). Current technology makes this much harder than it should be. Unit‑testing frameworks typically discard most observations (observable outputs to calls), preserving only pass/fail outcomes, and burying stimuli in scaffolding code. Execution tracing captures the opposite extreme – massive, low‑level call/state data – from which functional responses must be mined laboriously. In short, testing captures too little, while (extensive) tracing captures too much, and both obscure the functionality that is important.
@@
In earlier work, we addressed the representation gap by introducing three data structures tailored to the observation and recording of run‑time functional behavior [kesselOS2023]: Sequence Sheets to encode the precise invocations and the responses they elicit, the Stimulus-Response Matrices (SRMs) to compare many implementations against many tests, and Stimulus-Response Cubes (SRCs) (formely SRHs) to analyze ensembles of SRMs across additional dimensions such as repetitions, environments, and metrics. In an SRM, rows are tests and columns are implementations of the same functional abstraction (i.e., coding problem). Each cell stores the actuation (operation, inputs, outputs) for that test-implementation pair. SRCs bundle related SRMs into a multi‑dimensional structure for navigation and analysis. These structures make behavioral data analyzable offline through data-driven analytics, and reusable (replication, reproduction, repurposing) [kesselOS2023].
@@
While the efficacy of these data structures has been demonstrated, also at large scales [kesselOS2023, kesselNextGen2023], their current implementation in the LASSO platform assumes semi-sanitized scenarios where the tests and implementations that populate SRCs are “curated” for specific test-driven software experiments, or for particular diffential GAI applications [kesselGAI2024]. At present, they are not supported by any infrastructure that can collect this information, at scale and at high speed, from mainstream development environments and execution platforms, and they were not developed to allow SRC data to live alongside all the other data that is generated during software execution. More specifically, SRCs currently lack an environment/infrastructure that i) keeps every actuation, not just pass/fail verdicts; (ii) evolves without friction as new tests, implementations, environments, measurements and context appear; and (iii) answers interactive queries that reconstruct SRC views and support higher-level analyses (e.g., behavioral clustering, n-version assessment) without a distributed cluster of machines, and (iv) allows SRC data to live alongside other forms of structured and unstructured data within a broader record of run-time software execution. In short, we need an infrastructure that realizes the vision of continual SRCs – motivated recently in our research agenda for behavior-aware LLMs [kessel2025] – that grow as code is tested and make it integrable in and accessible from the current software engineering data tool landscape.
@@
To this end, we introduce the notion of an Observation Lakehouse that operationalizes the continual SRCs concept. Realized as a Python application based on a modern data lakehouse architecture [armbrust2021lakehouse], it uses Apache Iceberg [apache_iceberg] as a transactional table format over columnar Apache Parquet files [apache_parquet_format], with DuckDB for high-performance querying [duckdb]. Our approach unifies behavioral observations from two complementary streams – controlled software experiments and real-world CI pipelines – into a single, append-only data model. This foundation makes large-scale differential analysis practical and efficient, allowing users to perform n-version assessment of code variants, use behavioral clustering to identify equivalence classes, and compute consensus oracles without re-executing historical runs. The feasibility of this approach is demonstrated on a benchmark of over 8 million observation rows, where complex analytical queries – using SQL – complete in under 100ms on a standard laptop with a storage footprint of less than 51MiB. This paper contributes a lakehouse infrastructure that provides a practical foundation for mining software behavior at scale, addressing a key obstacle to creating more behavior-aware GAI.
@@
The remainder of the paper is organized as follows. Section II details the lakehouse architecture and the data-modeling choices that enable a continual SRC. Section III reports ingestion throughput, query latency, and the cost of behavioral clustering that underpins consensus‑oracle construction. Section IV discusses extensions (semantic search, Model‑Context‑Protocol server, federation) and outlines current limitations, before concluding with a call to the research community to adopt and extend the observation lakehouse.
@@ II Observation Lakehouse Architecture
In this section, we describe the design and implementation of the Observation Lakehouse, a platform that materializes the continual SRC [kessel2025]. The architecture is designed to satisfy four core requirements: (1) support for concurrent, incremental ingestion of both experimental and practitioner data; (2) unrestricted schema evolution for dimensional extensibility; (3) efficient on-the-fly reconstruction of arbitrary SRMs; and (4) fast, expressive analytics over massive observation sets.
@@ II-B Architectural Foundation: A Data Lakehouse
A conventional data lake (raw columnar files) lacks transactional guarantees, while a traditional RDBMS has a rigid schema and does not scale cost-effectively. To solve this, we employ a true data lakehouse architecture [armbrust2021lakehouse] that combines the strengths of both. Our stack consists of —
  • Columnar Storage (Apache Parquet [apache_parquet_format]): Provides efficient on-disk representation with high compression and predicate push-down (reduces how much data a query engine has to read from disk – avoiding full scans).
  • Transactional Table Format (Apache Iceberg [apache_iceberg]): Layered on top of Parquet, it provides ACID commits, snapshot isolation for time-travel, and, most importantly, schema evolution without rewriting existing files. This directly enables the SRC’s “dimensional extensibility”, allowing new columns to be added at any time.
  • SQL Analytics Engine (DuckDB [duckdb]): Offers a familiar, high-performance query interface for interactive analysis. As an efficient, in-process, and vectorized execution engine, its tight integration with Apache Arrow [apache_arrow] (an efficient in-Memory format) allows for zero-copy reads directly from Parquet files. This is ideal for providing near-instantaneous query feedback for analysts without the overhead of a distributed cluster.
@@ II-C Data Model and Physical Design
The logical SRC is materialized as three normalized Iceberg tables, mirroring a classic star-schema design known from data warehouses [kimball2013data]. This three-table model, detailed in Table I, separates the core entities of a test execution – the stimulus (i.e., test specification), the code unit under test, and the response. The primary advantage of this normalized design is massive de-duplication, as a single test or code implementation is stored only once.
@@
All three tables are physically partitioned on the two-level key (data_set_id, problem_id). This is the most crucial design choice for performance. It ensures that all data related to a single logical SRM is co-located on disk [duckdb_hive_partitioning]. When performing an analysis, the engine uses partition pruning to skip the vast majority of files and read only the relevant data, enabling efficient on-the-fly SRM reconstruction.
@@ II-D Illustrative Example: The Ingestion Flow
To make our data model concrete, consider the end-to-end ingestion flow for a stateful Queue implementation, an example used to demonstrate the capture of operational semantics in the original SRC concept [kessel2025]. First, a Java implementation of a queue is ingested into the code_implementations table and assigned a unique implementation_id (e.g., ’impl_queue_A’). Next, a test specification, with steps like create(), enqueue(1), and dequeue(), is ingested into the tests table and assigned a test_id (e.g., ’test_fifo_1’).
@@ II-F Analytical Capabilities Enabled by the Architecture
N-Version Assessment and Differential GAI [kesselGAI2024]: Users can run n-version assessments of many alternative implementations, including LLM-generated variants, under identical tests. Our platform facilitates this by enabling behavioral clustering to identify equivalence classes and compute consensus oracles (majority behavior) for regression checks.
@@ III Performance
The workload is derived from 509 coding problems (taken from [10103177]), a set of 13,384 LLM-generated implementations from four different LLMs, and a combination of 95,154 human-written and LLM-generated tests, yielding $\approx$8.6M step-level observations (see Table II). These observations were originally obtained using the LASSO platform [lasso_arena], a prototype system featuring an “arena” test driver for mass-execution. All measurements were performed on a laptop with an Intel Core i9, 64 GiB of RAM, and an NVMe SSD, running Ubuntu 24.04 and Python 3.12.
@@
Total Size 69.4iMB $\approx$ 8 MiB/1 million observations
The invocations in a call sequence include all invocations part of the entire test fixture (i.e., focal methods, test data initialization etc.). The high compression ratio – 8.6M step records fitting into $\approx$51MB (Table II) – is a direct consequence of the columnar Parquet format and its efficient run-length encoding of repeated values, such as implementation_id, within our normalized data model.
@@ III-A Ingestion Performance
Ingestion is performed in two phases: a one-time bulk import of all the code implementations and tests, which completes in $\approx$7 seconds, and the continuous ingestion of the observation records. We simulated the observation stream by writing the 8.6M records in batches using a Python worker with DuckDB and PyArrow. The total time to ingest all observations (including transformation costs) is $\approx$55 seconds, achieving a sustained ingestion rate of $\approx$155,000 records per second, confirming the feasibility of populating the lakehouse from high-throughput sources like CI pipelines.
@@ IV Future Work, Limitations, and Conclusion
Stimulus/response serialization is non-trivial for robust equivalence checking, especially for complex or non-JSON-friendly objects like binary streams. Our current JSON-based approach (cf. [lasso_arena]) supports polyglot interoperability, but future work will need to evolve this protocol, likely by incorporating content-based hashing or specialized handlers to compare syntactically-different values (e.g., exceptions) as equivalent outputs. This necessitates the use of custom equivalence functions – which could be realized as SQL User-Defined Functions.
@@
The implementation of our Observation Lakehouse is available as open source on GitHub [^1] https://github.com/SoftwareObservatorium/observation-lakehouse. The repository includes the full source code, instructions, dataset and interactive notebooks to replicate the performance evaluation.

### v2

@@ I Introduction
Most code-generating LLMs are trained on static artifacts – source code, comments, and specifications – rather than on materializations of run‑time behavior [6]. Such models inevitably internalize patterns from buggy or mislabeled code as readily as from correct implementations, because they are rarely, if ever, exposed to trustworthy evidence of functional run-time behavior (i.e., functionality) [8, 10]. Improving reliability, therefore, requires exposing models to ground‑truth behavior captured under execution.
@@
Rice’s theorem rules out deciding non‑trivial semantic properties of programs in general [19], so for real systems the only practical way of doing this is through dynamic observation – that is, to execute software where it is expected to meet its advertised functionality and record what it actually does, either during controlled tests (i.e., software testing [1]) or during real usage (i.e., monitoring and tracing). Current technology makes this much harder than it should be. Unit‑testing frameworks typically discard most observations (observable outputs to calls), preserving only pass/fail outcomes, and burying stimuli in scaffolding code. Execution tracing captures the opposite extreme – massive, low‑level call/state data – from which functional responses must be mined laboriously. In short, testing captures too little, while (extensive) tracing captures too much, and both obscure the functionality that is important.
@@
In earlier work, we addressed the representation gap by introducing three data structures tailored to the observation and recording of run‑time functional behavior [12]: Sequence Sheets to encode the precise invocations and the responses they elicit, the Stimulus-Response Matrices (SRMs) to compare many implementations against many tests, and Stimulus-Response Cubes (SRCs) (formerly SRHs) to analyze ensembles of SRMs across additional dimensions such as repetitions, environments, and metrics. In an SRM, rows are tests and columns are implementations of the same functional abstraction (i.e., coding problem). Each cell stores the actuation (operation, inputs, outputs) for that test-implementation pair. SRCs bundle related SRMs into a multi‑dimensional structure for navigation and analysis. These structures make behavioral data analyzable offline through data-driven analytics, and reusable (replication, reproduction, repurposing) [12].
@@
While the efficacy of these data structures has been demonstrated, also at large scales [12, 11], their current implementation in the LASSO platform assumes semi-sanitized scenarios where the tests and implementations that populate SRCs are “curated” for specific test-driven software experiments, or for particular differential GAI applications [14]. At present, they are not supported by any infrastructure that can collect this information, at scale and at high speed, from mainstream development environments and execution platforms, and they were not developed to allow SRC data to live alongside all the other data that is generated during software execution. More specifically, SRCs currently lack an environment/infrastructure that i) keeps every actuation, not just pass/fail verdicts; (ii) evolves without friction as new tests, implementations, environments, measurements and context appear; and (iii) answers interactive queries that reconstruct SRC views and support higher-level analyses (e.g., behavioral clustering, n-version assessment) without a distributed cluster of machines, and (iv) allows SRC data to live alongside other forms of structured and unstructured data within a broader record of run-time software execution. In short, we need an infrastructure that realizes the vision of continual SRCs – motivated recently in our research agenda for behavior-aware LLMs [13] – that grow as code is tested and make it integrable in and accessible from the current software engineering data tool landscape.
@@
To this end, we introduce the notion of an Observation Lakehouse that operationalizes the continual SRCs concept. Realized as a Python application based on a modern data lakehouse architecture [4], it uses Apache Iceberg [3] as a transactional table format over columnar Apache Parquet files [18], with DuckDB for high-performance querying [7]. Our approach unifies behavioral observations from two complementary streams – controlled software experiments and real-world CI pipelines – into a single, append-only data model. This foundation makes large-scale differential analysis practical and efficient, allowing users to perform n-version assessment of code variants, use behavioral clustering to identify equivalence classes, and compute consensus oracles without re-executing historical runs. The feasibility of this approach is demonstrated on a benchmark of over 8 million observation rows, where complex analytical queries – using SQL – complete in under 100ms on a standard laptop with a storage footprint of less than 51MiB. This paper contributes a lakehouse infrastructure that provides a practical foundation for mining software behavior at scale, addressing a key obstacle to creating more behavior-aware GAI.
@@
The remainder of the paper is organized as follows. Section II details the lakehouse architecture and the data-modeling choices that enable a continual SRC. Section III reports ingestion throughput, query latency, and the cost of behavioral clustering that underpins consensus‑oracle construction. Section IV discusses extensions (semantic search, MCP server, federation) and outlines current limitations, before concluding with a call to the research community to adopt and extend the observation lakehouse.
@@ II Observation Lakehouse Architecture
In this section, we describe the design and implementation of the Observation Lakehouse, a platform that materializes the continual SRC [13]. The architecture is designed to satisfy four core requirements: (1) support for concurrent, incremental ingestion of both experimental and practitioner data; (2) unrestricted schema evolution for dimensional extensibility; (3) efficient on-the-fly reconstruction of arbitrary SRMs; and (4) fast, expressive analytics over massive observation sets.
@@ II-B Architectural Foundation: A Data Lakehouse
A conventional data lake (raw columnar files) lacks transactional guarantees, while a traditional RDBMS has a rigid schema and does not scale cost-effectively. To solve this, we employ a true data lakehouse architecture [4] that combines the strengths of both. Our stack consists of —
  • Columnar Storage (Apache Parquet [18]): Provides efficient on-disk representation with high compression and predicate push-down (reduces how much data a query engine has to read from disk – avoiding full scans).
  • Transactional Table Format (Apache Iceberg [3]): Layered on top of Parquet, it provides ACID commits, snapshot isolation for time-travel, and, most importantly, schema evolution without rewriting existing files. This directly enables the SRC’s “dimensional extensibility”, allowing new columns to be added at any time.
  • SQL Analytics Engine (DuckDB [7]): Offers a familiar, high-performance query interface for interactive analysis. As an efficient, in-process, and vectorized execution engine, its tight integration with Apache Arrow [2] (an efficient in-Memory format) allows for zero-copy reads directly from Parquet files. This is ideal for providing near-instantaneous query feedback for analysts without the overhead of a distributed cluster.
@@ II-C Data Model and Physical Design
The logical SRC is materialized as three normalized Iceberg tables, mirroring a classic star-schema design known from data warehouses [16]. This three-table model, detailed in Table I, separates the core entities of a test execution – the stimulus (i.e., test specification), the code unit under test, and the response. The primary advantage of this normalized design is massive de-duplication, as a single test or code implementation is stored only once.
@@
All three tables are physically partitioned on the two-level key (data_set_id, problem_id). This is the most crucial design choice for performance. It ensures that all data related to a single logical SRM is co-located on disk [9]. When performing an analysis, the engine uses partition pruning to skip the vast majority of files and read only the relevant data, enabling efficient on-the-fly SRM reconstruction.
@@ II-D Illustrative Example: The Ingestion Flow
To make our data model concrete, consider the end-to-end ingestion flow for a stateful Queue implementation, an example used to demonstrate the capture of operational semantics in the original SRC concept [13]. First, a Java implementation of a queue is ingested into the code_implementations table and assigned a unique implementation_id (e.g., ’impl_queue_A’). Next, a test specification, with steps like create(), enqueue(1), and dequeue(), is ingested into the tests table and assigned a test_id (e.g., ’test_fifo_1’).
@@ II-F Analytical Capabilities Enabled by the Architecture
N-Version Assessment and Differential GAI [14]: Users can run n-version assessments of many alternative implementations, including LLM-generated variants, under identical tests. Our platform facilitates this by enabling behavioral clustering to identify equivalence classes and compute consensus oracles (majority behavior) for regression checks.
@@ III Performance
The workload is derived from 509 coding problems (taken from [5]), a set of 13,384 LLM-generated implementations from four different LLMs, and a combination of 95,154 human-written and LLM-generated tests, yielding $\approx$8.6M step-level observations (see Table II). These observations were originally obtained using the LASSO platform [17], a prototype system featuring an “arena” test driver for mass-execution. All measurements were performed on a laptop with an Intel Core i9, 64 GiB of RAM, and an NVMe SSD, running Ubuntu 24.04 and Python 3.12.
@@
Total Size 69.4MiB $\approx$ 8 MiB/1 million observations
The invocations in a call sequence include all invocations part of the entire test fixture (i.e., focal methods, test data initialization etc.). The high compression ratio – 8.6M step records fitting into $\approx$51MiB (Table II) – is a direct consequence of the columnar Parquet format and its efficient run-length encoding of repeated values, such as implementation_id, within our normalized data model.
@@ III-A Ingestion Performance
Ingestion is performed in two phases: a one-time bulk import of all the code implementations and tests, which completes in $\approx$7 seconds, and the continuous ingestion of the observation records. We simulated the observation stream by writing the 8.6M records in batches using a Python worker with DuckDB and PyArrow. The total time to ingest all observations (including transformation costs) is $\approx$55s, achieving a sustained ingestion rate of $\approx$155,000 records per second, confirming the feasibility of populating the lakehouse from high-throughput sources like CI pipelines.
@@ IV Future Work, Limitations, and Conclusion
Stimulus/response serialization is non-trivial for robust equivalence checking, especially for complex or non-JSON-friendly objects like binary streams. Our current JSON-based approach (cf. [17]) supports polyglot interoperability, but future work will need to evolve this protocol, likely by incorporating content-based hashing or specialized handlers to compare syntactically-different values (e.g., exceptions) as equivalent outputs. This necessitates the use of custom equivalence functions – which could be realized as SQL User-Defined Functions.
@@
The implementation of our Observation Lakehouse is available as open source on GitHub. The repository includes the full source code, instructions and interactive notebooks to replicate the performance evaluation. The dataset used for our performance evaluation is available in [15].